OpenStudy (anonymous):

Determine the volume of a 0.500M KMnO4 solution required to completely react with 2.85g of Zn

OpenStudy (paxpolaris):

\[2KMnO_4+Zn \rightarrow Zn \left( MnO_4 \right)_2+2K\]For this partial reaction you need twice as many moles of KMnO_4 as of Zn

OpenStudy (paxpolaris):

\[\large 2.85 g\ Zn \times{1mol\ Zn \over 65.38 g\ Zn}\times{2mol\ KMnO_4 \over 1mol Zn}\times{1\ liter\ solution \over 0.500 mol\ KMnO_4}=\]

OpenStudy (paxpolaris):

\[=2.85 \cancel {g\ Zn} \times{\cancel {1mol\ Zn }\over 65.38 \cancel {g\ Zn}}\times{2\cancel {mol\ KMnO_4} \over \cancel {1mol\ Zn}}\times{1\ liter\ solution \over 0.500 \cancel {mol\ KMnO_4}}\] \[=2.85*2/(65.38*.500)\ l=\Large 0.174\ liters\ of\ solution\]

OpenStudy (anonymous):

but the answer is wrong~

OpenStudy (paxpolaris):

what's the answer supposed to be ... maybe I didn't understand the reaction

OpenStudy (anonymous):

sorry..i dont know =)

OpenStudy (paxpolaris):

then how do you know it's wrong

OpenStudy (anonymous):

coz i have key in the answer and it says that my answer is wrong~

OpenStudy (paxpolaris):

idk ... maybe try 174ml or 175ml

OpenStudy (xishem):

Yep. The answer is definitely 174mL or 0.174L.